OpenStudy (tiffany_rhodes):

Evaluate the integral:

OpenStudy (dan815):

give it to me

OpenStudy (tiffany_rhodes):

\[\int\limits_{0}^{\pi} 12t^2\cos(t^3)dt\]

OpenStudy (luigi0210):

No me

OpenStudy (tiffany_rhodes):

I used u-sub and I know how to find the answer, i'm just confused on the endpoints of integration

OpenStudy (tiffany_rhodes):

I let u=t^3

OpenStudy (dan815):

ok yes

OpenStudy (tiffany_rhodes):

my final answer was 4sin(t^3)

OpenStudy (tiffany_rhodes):

I just need help on the endpoints :)

OpenStudy (luigi0210):

The upper and lower limits?

OpenStudy (tiffany_rhodes):

yes 0 to pi

OpenStudy (dan815):